OpenStudy (anonymous):

please help to solve the inverse of Y =X^3 -X +6

OpenStudy (anonymous):

\[ x = 1/6\,\sqrt [3]{-648+108\,y+12\,\sqrt {2904-972\,y+81\,{y}^{2}}}+\] \[ 2\,{ \frac {1}{\sqrt [3]{-648+108\,y+12\,\sqrt {2904-972\,y+81\,{y}^{2}}}}} \]